OpenStudy (anonymous):

Let u = x^2+ 3. Which equation is equivalent to (x^2+3)^2+ 21 = 10x^2+ 30 in terms of u?

OpenStudy (campbell_st):

well factor the right hand side of the equation and you get \[(x^2 + 3)^2 + 21 = 10(x^2 + 3)\] since \[u = x^2 + 3\] the equation becomes \[(u)^2 + 21 = 10(u)\] now just simplify

OpenStudy (anonymous):

Wait I typed in the equation incorrectly. I'm sorry. It was: Let u = x^2+ 3. (x^2+3)^2+ 21 = 10(x^2+ 30)

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Wait no. I'm sorry give me a second

OpenStudy (anonymous):

Sorry! it was originally correct, but what happened to the 30?

OpenStudy (campbell_st):

ok then so the left side is straight forward rewrite \[u = x^2 + 3~~~as~~~ x^2 = u - 3\] so the right side becomes \[10(x^2 + 30) = 10([u^2 - 3] + 30)\] hope that helps

OpenStudy (campbell_st):

oops should read \[10([u - 3] + 30)\]

OpenStudy (anonymous):

Okay I get it now! And the answer is u^2-10u+21?

OpenStudy (campbell_st):

well you have on the right side 10[u -3] + 30) = 10(u + 27) = 10u + 270 you can check to make sure

OpenStudy (anonymous):

woah.. you lost me now

OpenStudy (campbell_st):

the problem becomes \[u^2 + 21 = 10([u - 3] + 30) ~~~or~~~~u^2 + 21 = 10(u + 27)\] so distribute and collect like terms.